consider 교 73.0 kg man standing on spring sca e in an elevator, Starting rom rest the elavator ascends attaining its maximum speed o 1 the negative ydirection tor 1 s and comes to rest 6 m/s n a 550 s It travels with this constant spced or the nex t .00 s The clevator then undorgoes uniform acco cration n fe Whal dues the speing scele regiser before Le elevaur stars o muve? h) What doas the spring scala registar during tha first 1.550 s? (c) What does the spring scale register while the elevator is traveling at constant peed? (d) What does the spring scale register during the time it is slowing down?
Answer #1

a) F = ma = 73.0 * 9.81 = 716.13 N

b) acceleration a = dv / dt = 1.66 / 0.550

a = 3.02 m/s2

F = F' + m a = 716.13 + (73 * 3.02)

= 936.4 N

c) F = ma = 73.0 * 9.81 = 716.13 N

d) acceleration = dv/dt = -1.66 / 1.60 = -1.04 m/s2

F = F' + m a = 716.13 + (73 * -1.04) = 640 N